Solve. Identify all double roots.
(x-3)^3 - (x-3)^2 = 0

Answer by mananth(16946)

x-3)^3 - (x-3)^2 = 0
(x-3)^2(x-3-1)=0
(x-3)^2(x-2)=0
x=3 OR 2